# RobustFormer: Noise-Robust Pre-training for Images and Videos

### v1

@@ Abstract
While deep learning models are powerful tools that revolutionized many areas, they are also vulnerable to noise as they rely heavily on learning patterns and features from the exact details of the clean data. Transformers, which have become the backbone of modern vision models, are no exception. Current Discrete Wavelet Transforms (DWT) based methods do not benefit from masked autoencoder (MAE) pre-training since the inverse DWT (iDWT) introduced in these approaches is computationally inefficient and lacks compatibility with video inputs in transformer architectures.
In this work, we present RobustFormer, a method that overcomes these limitations by enabling noise-robust pre-training for both images and videos; improving the efficiency of DWT-based methods by removing the need for computationally iDWT steps and simplifying the attention mechanism. To our knowledge, the proposed method is the first DWT-based method compatible with video inputs and masked pre-training. Our experiments show that MAE-based pre-training allows us to bypass the iDWT step, greatly reducing computation. Through extensive tests on benchmark datasets, RobustFormer achieves state-of-the-art results for both image and video tasks.
Keywords Noise Robustnesss, Discrete Wavelet Transform, Masked Pre-training
@@ 1 Introduction
Addressing robustness in deep learning models for video data is crucial for practical applications, including surveillance, autonomous driving, and multimedia content analysis [34]. Just as images can suffer from distortions such as blurring, noise, and artifacts, video data experiences similar issues, compounded by the additional variability introduced over time [38]. Deep learning models, especially those trained on clean, high-quality data, are highly sensitive to these disruptions because they rely heavily on precise patterns within the data to make accurate predictions [48]. Even minor noise or inconsistencies in a single frame can cause the model to misinterpret important features, as these models have not been trained to ignore unexpected distortions [42]. With video data, the challenge becomes even greater due to the temporal nature of the information, where each frame builds on previous ones to create a coherent sequence [54]. For instance, temporal inconsistencies such as random noise across frames, motion blur, flickering shadows, and heat-induced turbulence can lead to abrupt changes between frames, confusing the model as it attempts to track and interpret motion or recognize actions [15, 54]. In tasks like autonomous driving, these inconsistencies are highly prevalent and pose significant risks, as the model may fail to accurately detect objects, assess distances, or predict movement patterns [40].
@@
A primary reason for these issues lies in the pixel-based design of most deep-learning models. Although RGB pixel representations are the standard in vision applications because they capture detailed color and spatial information, they are computationally demanding and highly sensitive to noise and domain shifts. This sensitivity arises since RGB data represents exact pixel-level details, so even slight distortions can disrupt deep-learning models that depend on these patterns for accuracy. Additionally, RGB representations process both essential and irrelevant features equally, limiting their efficiency. Inspired by image compression techniques like JPEG, frequency domain transformations such as Discrete Cosine Transform (DCT) provide a solution by isolating low-frequency components (representing significant information) from high-frequency ones (less relevant details) [32, 39]. This separation reduces memory demands and increases robustness to noise without compromising performance [11, 16, 32].
@@
While pure frequency domain transformations offer these benefits, they also lead to a loss of spatial or temporal locality due to the change in basis, a consequence of the Heisenberg uncertainty principle. In contrast, Discrete Wavelet Transform (DWT) provides a more balanced approach, decomposing the signal into both high- and low-frequency components across multiple scales. This multi-resolution representation allows DWT to retain spatial and temporal information, preserving important features and contextual details while still isolating high-frequency information. Image-based methods leveraging DWT have shown promise in enhancing robustness against spatial noise by allowing models to focus on frequency-based features that are less affected by disruptions [30]. Applying these principles to video data can address both temporal and spatial corruptions, as DWT can help models focus on low-frequency information, improving robustness to noise and other distortions across frames.
@@
In this paper, we employ a Video Masked Autoencoding approach [49, 50], additionally, incorporating DWT for pre-training and fine-tuning on the uncorrupted datasets, followed by evaluation on their corrupted counterparts. We use UCF101 and HMDB51 datasets for training and evaluation, where the generated corruptions are of 6 types(including 5 spatial and 1 temporal corruptions in 5 severity levels) reflecting real world scenarios [57].
@@
1.
We present a novel Discrete Wavelet Transform (DWT) based masked autoencoder architecture that is robust to spatial and temporal corruptions in both video and image data. To our knowledge this is the first work to implement DWT in Masked autoencoder setting.
2.
We perform comprehensive evaluation on several real-world noise types with varying severity levels in large-scale benchmark datasets.
3.
We demonstrate that our method performs equally well and in many cases better than commonly used IDWT variant which requires comparatively more compute resource and makes architectures overly complicated.
@@ 2.1 Masked Video Training
BERT introduced the concept of Masked Language Modeling (MLM), a novel pretraining objective for natural language understanding, which led to significant advances in NLP [26]. Following BERT’s success, various NLP models, including RoBERTa [35], ALBERT [28], and ELECTRA [6], extended this masked training approach. Inspired by BERT’s masked token prediction, the concept of Masked Image Modeling (MIM) emerged for computer vision tasks. iGPT [5] first adapted the Transformer model for image data by treating images as pixel sequences and using a similar masked prediction task. BEiT [1] extended this concept by proposing a discrete variational autoencoder (dVAE) to tokenize image patches, predicting masked tokens in a BERT-style pretraining. These approaches demonstrated that MIM could effectively learn transferable representations for various downstream tasks. Masked Autoencoders (MAE) [19] refined MIM by reconstructing masked regions in images, achieving state-of-the-art results on multiple benchmarks using Vision Transformers (ViTs) and large-scale unlabeled data.
@@ 2.2 Discrete Wavelet transform(DWT)
Wavelets are essential in time-frequency analysis and signal processing tasks, such as anti-aliasing and detail restoration, through Discrete Wavelet Transform (DWT) and its inverse (IDWT) [36]. Early studies combined wavelets with shallow neural networks for function approximation and classification, optimizing wavelet parameters within the network [47]. Recently, deeper networks have adopted wavelets for image classification, although these integrations can be computationally intensive [7]. The Multi-level Wavelet CNN (MWCNN) [33] applies Wavelet Packet Transform (WPT) for image restoration, handling both low- and high-frequency components. Similarly, the Convolutional-Wavelet Neural Network (CWNN) [10] utilizes dual-tree complex wavelet transform (DT-CWT) to reduce noise in SAR images while preserving crucial features, though within a simplified two-layer structure. Wavelet Pooling [53] employs a two-level DWT for pooling and combines DWT/IDWT for backpropagation, deviating from conventional gradient methods.
A key advantage of wavelets is their ability to provide sparse representations, which enables more efficient processing and faster model training [53, 31, 54]. DWT-based training reduces redundancy, making it well-suited for real-time video classification tasks where small shifts in data can significantly affect model output. Unlike other approaches, our method avoids reconstruction after the wavelet transform, thus saving considerable computational resources, especially when handling high-resolution video data. This streamlined process enhances efficiency while maintaining robust performance in noisy environments.
@@ 2.3 Noise Robustness
Since DWT has been widely applied in noisy data reconstruction, achieving notable gains in both noise removal and complex tasks like deblurring [53, 31], our approach builds on the methodologies of WaveFormer [55] and WaveCNets [30] by incorporating DWT. As illustrated in Figure 2, (b) shows WaveFormer’s framework, which computes attention using only the low-frequency components of the query and key matrices. Unlike WaveFormer, however, we omit DWT and inverse DWT (IDWT) for the value matrix, passing it directly. Similar to WaveCNets, we discard high-frequency components during the initial downsampling layer but extend this technique to 3D-DWT to handle temporal decomposition.
@@ 3 Methods
Compared to the classical VideoMAE [48], our proposed method enhances robustness by employing 3D-DWT for spatio-temporal analysis; in contrast to WaveFormer [54], our approach eliminates the computationally expensive iDWT step, significantly improving efficiency, as illustrated in Figure 2.
Our implementation incorporates a two-step wavelet transform. First, we utilize DWT’s efficient downsampling capability to generate embeddings for video patches. By leveraging 3D-DWT, we can process video data more effectively than stacking coefficients from 2D-DWT. This approach enables simultaneous handling of both temporal and spatial noise, ensuring robust feature extraction. The forward and backward propagation processes for this step are detailed below.
@@ 3.1 Wavelet Embedding
Let be the original downsampled and clipped video sequence of sequence length , height , width and channels . We start with the generation of transformation matrices based on the selected wavelet filters. This step is critical as it defines the low-pass and high-pass filters that are applied in the transformation process. Wavelet Filter Coefficients: Define the wavelet by its low-pass (L) and high-pass (H) filter coefficients.
Now we construct the transformation matrices for each dimension (depth, height, and width). For simplicity, only the depth dimension’s matrix construction is shown here. We can replicate this for the height and width dimensions:
@@ 3.2 Attention Computation
In the wavelet embedding step we average, concat or omit the high frequency components in an effort to avoid noise and smooth out the input representation. However since we perform additional latent transformation(3D Convolution with tube masking) which might escalate some of the remaining noises during attention calculation. We want to minimize the redundant details as much as possible during our overall training. Prior works[54] have proved that noise before the attention mechanism alters the correlation scores between features by increasing the values for unrelated pairs (due to the random alignment by noise) and decreasing them for related pairs (as noise dilutes the true alignment), thereby distorting the focus of the attention mechanism. This in our case leads to highly deviated classification accuracy. We thus follow steps from [54] to smooth out the latent representation before proceeding to attention calculation by completely omitting the involved high frequency components.
Suppose denotes the output of the tube-masked 3D convolution, representing the deep encoded feature representations. Here, each is a tensor in , where is the tubelet size, is the embedding size, and is the patch size. The elements of the attention mechanism, , , and , are first calculated by passing through three different linear layers.
We then perform 1D-DWT for each Q, K and V by obtaining the low and high pass filters corresponding to a specific wavelet(for eg. haar) similar to [54] as follows,
We calculate these segregated components similar to 2 and omit and . Note we also omit dwt computation for , so that the attention computation can attend to some useful high level components from (that are omitted in and in the input, as shown in Figure 3. With this we can avoid performing additional IDWT step as in [54] which makes the architecture complicated and computationally intensive.
We finally perform our attention score computation as follows:
@@
With this we create various configurations of Robustformer that involves various operations over DWT blocks. RF-A indicates Robustformer with dwt averaging, RF-AA refers to averaging as well as dwt attention as mentioned in 5. Similarly we define RF-O, RF-OA, RF-C and RF-CA for omit, omit using dwt attention, concat and concat using dwt attention.
@@
AlexNet | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 97 | 98 | 96 | 109 | 134 |
SqueezeNet | 107 | 106 | 105 | 103 | 101 | 100 | 100 | 101 | 103 | 97 | 98 | 98 | 100 | 98 | 106 |
VGG-11 | 97 | 97 | 100 | 92 | 93 | 91 | 94 | 94 | 94 | 89 | 84 | 75 | 84 | 78 | 87 |
VGG-19 | 92 | 89 | 91 | 89 | 89 | 89 | 90 | 89 | 89 | 86 | 78 | 64 | 80 | 84 | 85 |
VGG-19+BN | 82 | 83 | 88 | 84 | 84 | 83 | 89 | 82 | 87 | 85 | 68 | 59 | 87 | 74 | 85 |
ResNet-18 | 87 | 88 | 91 | 84 | 91 | 87 | 88 | 91 | 88 | 78 | 69 | 69 | 84 | 88 | 80 |
ResNet-50 | 80 | 82 | 83 | 75 | 89 | 78 | 80 | 78 | 78 | 75 | 66 | 57 | 71 | 85 | 87 |
@@ 4.1 Datasets
We evaluate our method across five datasets: Kinetics-400 [25] and UCF-101 [46] for video tasks, and ImageNet-1K [8], ImageNet-Tiny-200 [29], and HMDB-51 for image tasks.
The Kinetics-400 dataset includes 240k training videos and 20k validation videos, each clip lasting 10 seconds. The dataset focuses on human-centered actions, covering a wide range of interactions. The UCF-101 dataset, while smaller with 13.3k action videos across 101 categories, offers high diversity in actions and significant variations in camera motion, object appearance, pose, background, and lighting conditions. For image-based experiments, we use the standard ImageNet-1K dataset. Additionally, we include ImageNet-Tiny-200 and the smaller video variant HMDB-51 to demonstrate our method’s effectiveness on smaller datasets, where Vision Transformers (ViTs) often struggle.
For noisy evaluation on video datasets, we assess our models under real-world perturbations following [42], along with additional noise types and intensity configurations. Video noise types include four categories: Noise (Gaussian, shot, impulse, speckle), Blur (zoom, motion, defocus), Digital (JPEG and MPEG compression artifacts), Temporal (jumble, box jumble), and Camera Motion (static rotation, dynamic rotation, translation). Each noise type is tested across five severity levels, resulting in a total of 70 unique noise configurations. In the image experiments, we evaluate model robustness using the ImageNet-P and ImageNet-C benchmarks [23]. ImageNet-C includes 15 types of corruptions, each with five severity levels, totaling 75 distinct corruptions. ImageNet-P adds 10 perturbation types with 30 intensity variations, leading to 300 different noise scenarios. All pre-training, fine-tuning, and validation are conducted on clean datasets, without noise augmentation.

### v2

@@ Abstract
While deep learning-based models like transformers, have revolutionized time-series and vision tasks, they remain highly susceptible to noise and often overfit on noisy patterns rather than robust features. This issue is exacerbated in vision transformers, which rely on pixel-level details that can easily be corrupt. To address this, we leverage the discrete wavelet transform (DWT) for its ability to decompose into multi-resolution layers, isolating noise primarily in the high frequency domain while preserving essential low-frequency information for resilient feature learning. Conventional DWT-based methods, however, struggle with computational inefficiencies due to the requirement for a subsequent inverse discrete wavelet transform (IDWT) step. In this work, we introduce RobustFormer, a novel framework that enables noise-robust masked autoencoder (MAE) pre-training for both images and videos by using DWT for efficient downsampling, eliminating the need for expensive IDWT reconstruction and simplifying the attention mechanism to focus on noise-resilient multi-scale representations. To our knowledge, RobustFormer is the first DWT-based method fully compatible with video inputs and MAE-style pre-training. Extensive experiments on noisy image and video datasets demonstrate that our approach achieves up to 8% increase in Top-1 classification accuracy under severe noise conditions in Imagenet-C and up to 2.7% in Imagenet-P standard benchmarks compared to the baseline and up to 13% higher Top-1 accuracy on UCF-101 under severe custom noise perturbations while maintaining similar accuracy scores for clean datasets. We also observe the reduction of computation complexity by up to 4.4% through IDWT removal compared to VideoMAE baseline without any performance drop.
@@ 1 Introduction
Addressing robustness in deep learning models for images and video data is crucial for practical applications, including surveillance, autonomous driving, and multimedia content analysis [34]. Images can suffer from distortions such as blurring, noise, and artifacts, while video data experience similar issues, compounded by the additional variability introduced over time [38]. Deep learning models, especially those trained on clean, high-quality data, are highly sensitive to these disruptions because they rely heavily on precise patterns within the data to make accurate predictions [50]. Even minor noise or inconsistencies in an image can cause the model to misinterpret important features, as these models are not inherently robust to unexpected distortions [43]. With video data, the challenge becomes even greater due to the temporal nature of the information, where each frame builds upon previous ones to create a coherent sequence [55]. For instance, temporal inconsistencies such as random noise across frames, motion blur, flickering shadows, and heat-induced turbulence can lead to abrupt changes between frames, confusing the model as it attempts to track and interpret motion or recognize actions [14, 55]. In tasks like autonomous driving, these inconsistencies are highly prevalent and pose significant risks, as the model may fail to accurately detect objects, assess distances, or predict movement patterns [41].
@@
A primary reason for these issues lies in the pixel-based design of most deep-learning models. Although RGB pixel representations are the standard in vision applications because they capture detailed color and spatial information, they are computationally demanding and highly sensitive to noise and domain shifts. This sensitivity arises since RGB data represent exact pixel-level details, so even slight distortions can disrupt deep learning models that depend on these patterns for accuracy. Additionally, RGB representations process both essential and irrelevant features equally, limiting their efficiency. Inspired by image compression techniques like JPEG, frequency domain transformations such as Discrete Cosine Transform (DCT) provide a solution by isolating low-frequency components (representing significant information) from high-frequency ones (less relevant details) [32, 39]. This separation reduces memory demands and increases noise resistance without compromising performance [11, 16, 32].
@@
While pure frequency domain transformations offer these benefits, they also lead to a loss of spatial or temporal locality due to the change in basis, a consequence of the Heisenberg uncertainty principle. In contrast, Discrete Wavelet Transform (DWT) provides a more balanced approach, decomposing the signal into both high- and low-frequency components across multiple scales. This multi-resolution representation allows DWT to retain spatial and temporal information, preserving important features and contextual details while still isolating high-frequency information. Image-based methods that leverage DWT have shown promise in improving robustness against spatial noise by allowing models to focus on frequency-based features that are less affected by disturbances [30]. Applying these principles to images and videoscan address both spatial and temporal (for videos) corruptions, as DWT can help models focus on low-frequency information, improving robustness to noise and other distortions across frames.
@@
In this paper, we employ a masked autoencoding approach [49, 51], additionally, incorporating DWT for pre-training and fine-tuning on the uncorrupted datasets, followed by evaluation on their corrupted counterparts. We experiment on widely used datasets for both images and videos, where the generated corruptions include both spatial and temporal types reflecting real-world scenarios [57].
@@
We present a novel Discrete Wavelet Transform (DWT) based masked autoencoder architecture that is robust to spatial and temporal corruptions in both video and image data. To our knowledge, this is the first work to implement DWT in a masked autoencoder setting.
We perform a comprehensive evaluation of several real-world noise types with varying severity levels in large-scale benchmark datasets.
We demonstrate that our method performs equally well and in many cases better than the commonly used IDWT variant, which requires comparatively more compute resource and makes architectures overly complicated.
@@ 2.1 Masked Video Training
BERT introduced the concept of masked language modeling (MLM), a novel pre-training objective for natural language understanding, which led to significant advances in NLP [26] and further inspiring extensions including RoBERTa [35], ALBERT [28], and ELECTRA [6], extended this masked training approach. Also inspired by BERT’s masked token prediction, the concept of masked image modeling (MIM) emerged for computer vision tasks. iGPT [5] first adapted the Transformer model for image data by treating images as pixel sequences and using a similar masked prediction task. BEiT [1] extended this concept by proposing a discrete variational autoencoder (dVAE) to tokenize image patches, predicting masked tokens in a BERT-style pre-training. These approaches demonstrated that MIM could effectively learn transferable representations for various downstream tasks. Masked Autoencoders (MAE) [19] refined MIM by reconstructing masked regions in images, achieving state-of-the-art results on multiple benchmarks using Vision Transformers (ViTs) and large-scale unlabeled data.
@@ 2.2 Discrete Wavelet transform(DWT)
Wavelets are essential in time-frequency analysis and signal processing tasks, such as anti-aliasing and detail restoration, through Discrete Wavelet Transform (DWT) and its inverse (IDWT) [36]. Early studies combined wavelets with shallow neural networks for function approximation and classification, optimizing wavelet parameters within the network [48]. Recently, deeper networks have adopted wavelets for image classification, although these integrations can be computationally intensive [7]. The Multilevel Wavelet CNN (MWCNN) [33] applies Wavelet Packet Transform (WPT) for image restoration, handling both low- and high-frequency components. Similarly, the Convolutional-Wavelet Neural Network (CWNN) [10] utilizes dual-tree complex wavelet transform to reduce noise in SAR images while preserving crucial features, though within a simplified two-layer structure. Wavelet Pooling [54] employs a two-level DWT for pooling and combines DWT/IDWT for backpropagation, deviating from conventional gradient methods.
A key advantage of wavelets is their ability to provide sparse representations, which enables more efficient processing and faster model training [54, 31, 55]. DWT-based training reduces redundancy, making it well-suited for real-time video classification tasks where small shifts in data can significantly affect model output. Unlike other approaches, our method avoids reconstruction after the wavelet transform, thus saving considerable computational resources, especially when handling images and videos. This streamlined process enhances efficiency while maintaining robust performance in noisy environments.
@@ 2.3 Noise Robustness
Since DWT has been widely applied in noisy data reconstruction, achieving notable gains in both noise removal and complex tasks like deblurring [54, 31], our approach builds on the methodologies of WaveCNets [30] by incorporating DWT for noise-robust representation. Figure 2(b) shows IDWT configuration which computes attention using only low-frequency components of the query and key for noise filtering(similar to [55]) during attention computation. Similar to WaveCNets, we discard high-frequency components during the initial downsampling layer but extend this technique to 3D-DWT to handle temporal decomposition. Our framework leverages 3D-DWT’s ability to decompose both spatially and temporally, enabling effective noise handling across frames and enhancing robustness over the entire video sequence. As shown in Figure 1, our approach improves both accuracy and relative robustness compared to several action recognition models on the UCF-101 dataset.
@@ 3 Methods
Compared to the classical VideoMAE [50], our proposed method enhances robustness by employing 3D-DWT for spatio-temporal analysis; in contrast to other IDWT-based methods, our approach eliminates the computationally expensive IDWT step, significantly improving efficiency, as illustrated in Figure 2(c).
Our implementation incorporates a two-step wavelet transform. First, we utilize DWT’s efficient downsampling capability to generate embeddings for video patches. By leveraging 3D-DWT, we can process video data more effectively than stacking coefficients from 2D-DWT. This approach enables simultaneous handling of both temporal and spatial noise. The robust feature extraction achieved is thus the result of noise-adaptive compression in the initial DWT phase, noise filtering during attention computation well as masked pretraining which allows to recover back to RGB without requiring additional IDWT step. The forward and backward propagation processes for this step are detailed below. For experiments with images we use 2D-DWT instead on 3-dimensional version while every other aspects remaining same.
@@ 3.1 Wavelet Embedding
Let be the original downsampled and clipped video sequence of sequence length , height , width and channels . We start with the generation of transformation matrices based on the selected wavelet filters. This step is critical as it defines the low-pass and high-pass filters that are applied in the transformation process. We define the wavelet by low-pass (L) and high-pass (H) filter coefficients as and . In this work, we employ Haar wavelets due to their computational efficiency and effectiveness in noise reduction.
Now we construct the transformation matrices for each dimension (depth, height, and width). For simplicity, only the depth (time) dimension’s matrix construction is shown here. We can replicate this for the height and width dimensions:
@@ 3.2 Attention Computation
In the wavelet embedding step, we average, concatenate, or omit high-frequency components to reduce noise and smooth the input representation. However, the subsequent latent transformation (3D convolution with tube masking) may amplify residual noise during attention calculation. To minimize redundant details, we aim to reduce noise as much as possible during training. Prior work [55] has shown that noise before the attention mechanism distorts correlation scores by increasing values for unrelated pairs (due to random alignment) and decreasing values for related pairs (as noise weakens true alignment), ultimately impairing classification accuracy. To mitigate this, we follow [55] and fully omit high-frequency components, ensuring a smoother latent representation before attention calculation.
Suppose denotes the output of the tube-masked 3D convolution, representing deep encoded feature representations. Each is a tensor in , where denotes the tubelet size, the embedding dimension, and the patch size. The attention mechanism elements , , and are computed by passing through three distinct linear layers.
We then perform 1D-DWT for each Q, K and V by obtaining the low and high pass filters corresponding to a specific wavelet(Haar) similar to [55] as follows,
We calculate these segregated components similar to eq (2) and omit and . Note that we also omit DWT computation for , so that the attention computation can attend to some useful high level components from (that are omitted in and in the input, as shown in Figure 3. With this we can avoid performing additional IDWT step as in [55] which makes the architecture complicated and computationally intensive.
We finally compute our attention score as follows:
@@
With this we create various configurations of RobustFormer that involves various operations over DWT blocks. RF-A indicates RobustFormer with DWT averaging, RF-AA refers to averaging as well as DWT attention to every attention layer as mentioned in 5. Similarly we define RF-O, RF-OA, RF-C and RF-CA for omit, omit using DWT attention, concat and concat using DWT attention.
@@
| AlexNet | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 |
| SqueezeNet | 107 | 106 | 105 | 100 | 103 | 101 | 100 | 101 | 103 | 97 | 97 | 98 | 106 | 109 | 134 |
| VGG-11 | 97 | 97 | 100 | 92 | 99 | 93 | 91 | 92 | 91 | 84 | 75 | 86 | 97 | 107 | 100 |
| VGG-19 | 89 | 91 | 95 | 89 | 98 | 90 | 90 | 89 | 86 | 75 | 68 | 80 | 97 | 102 | 94 |
| VGG-19+BN | 82 | 83 | 88 | 82 | 94 | 84 | 86 | 80 | 78 | 69 | 61 | 74 | 94 | 85 | 83 |
| ResNet-18 | 87 | 88 | 91 | 84 | 91 | 87 | 89 | 86 | 84 | 78 | 69 | 78 | 90 | 80 | 85 |
| ResNet-50 | 80 | 82 | 83 | 75 | 89 | 78 | 80 | 78 | 75 | 66 | 57 | 71 | 85 | 77 | 77 |
@@ 4.1 Datasets
We evaluate our method across five datasets: Kinetics-400 [25], UCF-101 [47] and HMDB-51 for video tasks, and ImageNet-1K [8] and ImageNet-Tiny-200 [29], for image tasks. The Kinetics-400 dataset includes 240k training videos and 20k validation videos, each clip lasting 10 seconds. The dataset focuses on human-centered actions, covering a wide range of interactions. The UCF-101 dataset, while smaller with 13.3k action videos across 101 categories, offers high diversity in actions and significant variations in camera motion, object appearance, pose, background, and lighting conditions. For image-based experiments, we use the standard ImageNet-1K dataset.
For noisy evaluation on video datasets, we assess our models under real-world perturbations following [43], along with additional noise types and intensity configurations. Video noise types include four categories: Noise (Gaussian, shot, impulse, speckle), Blur (zoom, motion, defocus), Digital (JPEG and MPEG compression artifacts), Temporal (jumble, box jumble), and Camera Motion (static rotation, dynamic rotation, translation). Each noise type is tested across five severity levels, resulting in a total of 70 unique noise configurations. In the image experiments, we evaluate model robustness using the ImageNet-P and ImageNet-C benchmarks [23]. ImageNet-C includes 15 types of corruptions with five severity levels, totaling 75 distinct corruptions. ImageNet-P adds 10 perturbation types with 30 intensity variations, leading to 300 different noise scenarios. All pre-training, fine-tuning, and validation are conducted on clean datasets, without noise augmentation.